OpenStudy (anonymous):

integrate x/1+x^4 dx

OpenStudy (anonymous):

x^5/(5) + x^2/(2) + C

OpenStudy (anonymous):

no the question is\[\int\limits x/1+x ^{4} dx\]

OpenStudy (anonymous):

yes and my answer corresponds to that

OpenStudy (anonymous):

so this integral can be interpreted by getting it in the form of 1/1+u^2 du. If we let u = x^2, then du = 2xdx do now the integral becomes\[(1/2)\int\limits_{}^{}1/(1+u^2)du\] this equals (1/2)arctan(u) + c, so your final answer is \[(1/2)\arctan(x^2)+C\]